When 1.00g of acetylene, C2H2, is burned, 48.2 kJ of heat is given off. When 0.750g of acetylene is burned in a bomb calorimeter (heat capacity = 1.117kJ/°C), the final temperature of the bomb calorimeter is 54.5°C. What is the initial temperature of the calorimeter when the reaction is complete?
1
Expert's answer
2019-12-16T04:53:23-0500

Mass=1g, Q= 48.2Kj

When mass,m=075g, Tf= 54.5°C, c=1.117kJ/°C

If 1g gives 48.2kj, then 0.75g will give 36.15Kj

"Q= m*c*dT"

Where "dT=Tf -Ti"

36.15=0.75*1.117*(54.5-Ti)

36.15=45.657375- 0.83775Ti

Ti=11.3°C